The volumes of soda in quart-sized soda bottles are normally distributed with a mean of 32.3 ounces anda standard deviation of 1.2 ounces. What is the probability that the volume of soda in a randomly selected bottle will be more than 31 ounces?
Answer #1

Solution :

Given that,

mean = = 32.3

standard deviation = = 1.2

P(x >31 ) = 1 - P(x< 31)

= 1 - P[(x -) / < (31-32.3) /1.2 ]

= 1 - P(z <-1.08 )

Using z table

= 1 -  0.1401

=0.8599